If tan thita + sec thita=4,then find sin thita. (0

Grade:11

2 Answers

Ashwin Muralidharan IIT Madras
12 years ago

Hi Arnab.

 

tanx + secx = 4

 

We also know that sec^2 x - tan^2 x = 1

or (secx-tanx)(secx+tanx) = 1

or secx-tanx = 1/4

 

So secx = 17/8 and tanx = 15/8

 

So sinx = tanx/secx = 15/17.

 

Hope that helps.

 

Best Regards,

Ashwin (IIT Madras).

sanatan sharma
12 years ago

the answer will be , sin Θ =  15/17 , -1